OpenStudy (anonymous):

Find the differential dy y=(squareroot 36-x^2)

OpenStudy (nottim):

eekkk

OpenStudy (anonymous):

Help :(

OpenStudy (anonymous):

let me rewrite this equation @Roro567, because I often start and then later on I figure out that I am not talking about the same equation as in the opening post \[ y= \sqrt{36-x^2}\] is that the equation?

OpenStudy (nottim):

sure is, spacey

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

then use the chain rule \[ \frac{dy}{dx}= \frac{-x}{\sqrt{36-x^2}} \]

OpenStudy (nottim):

yayy. might need to explain that more spacey

OpenStudy (anonymous):

You can split it up into two differential if you like, using substitution.

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

\[ u = 36-x^2 \] so \(y=\sqrt{u}\) then you can find both differentials and multiply them together. You know how to do that?

OpenStudy (anonymous):

yes thanks!

OpenStudy (anonymous):

very well.